Write the equation for the parabola that goes through the point (9,3) and has a vertex of (6,-5).
Thank You

Answer by lwsshak3(11628)

Write the equation for the parabola that goes through the point (9,3) and has a vertex of (6,-5).
**
Standard form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex.
For given parabola:
using given point and coordinates of vertex:
equation: 3=A(9-6)^2-5
3=9A-5
9A=8
A=8/9
Equation of given parabola:
y=(8/9)(x-6)^2-5